OpenStudy (anonymous):

The formula I = p r t gives the amount of simple interest (I) earned by investing principle ( p ) for rate ( r ) over a period of time ( t ). Solve this formula for p, and find p if r = 0.035, t = 5 and I = $350.

OpenStudy (anonymous):

\[I = p*r*t\] \[350 = p*0.035*5\] \[p=\frac{ 350 }{ 0,0035*5 }\] \[p=20000\]